OpenStudy (anonymous):

solve for x. cuberoot(x-2) -8=8

OpenStudy (anonymous):

\[\sqrt[3]{x-2} - 8 = 8\]

OpenStudy (anonymous):

yes, just like that ^

OpenStudy (anonymous):

x = 4098 Cube both sides. x-2 = 16^3 x = 4096 + 2 x = 4098

OpenStudy (anonymous):

ohh, alright! thanks

OpenStudy (anonymous):

\[(x-2)^\frac{1}{3}=4096^\frac{1}{3}=>x=4096+2=4098\]